# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ 2.4 The ALG as a stochastic process
So far, we have described the ALG that results when a realisation of the forward model is given. We now define it as a stochastic process independent of such a realisation, again first in the untyped version denoted by $(\Lambda(\tau))_{\tau\geq 0}$, and then in the typed version.
@@
    For the untyped version, let $\Lambda_{\alpha}(\tau)\subseteq[N]$ denote the set of ancestors at backward time $\tau$ of individual $\alpha\in[N]$ at backward time 00; and, for $G\subseteq[N]$, set $\Lambda_{G}(\tau)\mathrel{\mathop{:}}=\cup_{\alpha\in G}\Lambda_{\alpha}(\tau)$ by slight abuse of notation (so $\Lambda_{\{\alpha\}}(\tau)=\Lambda_{\alpha}(\tau)$). Then the process $(\Lambda(\tau))_{\tau\geq 0}\mathrel{\mathop{:}}=\big{(}\Lambda_{1}(\tau),\ldots,\Lambda_{N}(\tau)\big{)}_{\tau\geq 0}$ contains the complete genealogical information of all individuals in the present population. It evolves as follows. Clearly $(\Lambda_{1}(0),\ldots,\Lambda_{N}(0))=(\{1\},\{2\},\ldots,\{N\})$. If, for any time $\tau\geq 0$, the current state is $\Lambda(\tau)=(\lambda_{1},\ldots,\lambda_{N})$, the following events may happen: at rate $Nu$, a death event occurs and a randomly-chosen individual $\beta$ dies. So $\beta$ is removed from the set of ancestors of anybody, and
@@
    At rate $Ns$, a learning event occurs and a randomly-chosen individual $\beta_{1}$ learns from another randomly chosen individual $\beta_{2}$ (where $\beta_{2}=\beta_{1}$ is allowed, in which case the event is silent). So if, for $\alpha\in[N]$, $\beta_{1}$ is contained in $\lambda_{\alpha}$, then $\beta_{2}$ joins this ancestral set, and otherwise nothing happens:
    $\lambda_{\alpha}\longrightarrow\begin{cases}\lambda_{\alpha}\cup\{\beta_{2}\},&\text{if }\beta_{1}\in\lambda_{\alpha},\\ \lambda_{\alpha},&\text{otherwise}\end{cases}\qquad\alpha\in[N].$ (6b)
@@
    The untyped ALG may be turned into a typed one by superposing the untyped graph with a Poisson point process that lays down innovations at rate $\mu$ on every line. Together with an initial assignment of types to the lines alive at backward time $t$ and propagation along the graph as in the forward model, this determines the types along all lines in the graph.
Returning to the untyped version, let us note for later use that, since $\Lambda_{\alpha}(\tau)\subseteq\Lambda_{[N]}(\tau)$,
[…]
for any $\alpha\in[N]$.
In what follows, an important role will be played by the line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ of the ALG. That is,
$Y_{n}(\tau)\mathrel{\mathop{:}}=\lvert\Lambda_{[n]}(\tau)\rvert$ (8)
is the number of lines at backward time $\tau$ of the sample $[n]$ of individuals taken at backward time 00; due to exchangeability, we allow ourselves to simply speak of a sample of size $n$. Also, we will sometimes omit the dependence on the initial value $n$. $(Y(\tau))_{\tau\geq 0}$ is a birth-death process [^1] This birth-death process is not to be confused with the death-birth events in the model, as indicated by crosses in the graphical representation. on $[N]_{0}:=\{0,1,\ldots,N\}$ with birth and death rates222not to be confused with the realisations $\lambda_{\alpha}$ of $\Lambda_{\alpha}$ and the innovation rate $\mu$.
@@
respectively, when $Y(\tau)=y$ (note that the factor $(N-y)/N$ is the probability that a learning arrow comes from outside the current graph). Clearly, $(Y(\tau))_{\tau\geq 0}$ is an absorbing Markov chain with 0 as the only absorbing state. The behaviour of the process is very well studied, since it is, at the same time, the stochastic logistic process of ecology and epidemiology; in particular, $(Y(\tau))_{\tau\geq 0}$ is the number of infected individuals in a stochastic SIS model with infection rate $s$ and recovery rate $u$ per individual (Andersson and Britton, 2000, Ch. 8.2).
@@
Another crucial quantity will be $L_{n}$, the tree length, that is, the total length of all branches, as $\tau\to\infty$, in the ALG of a sample of size $n$, that is,
@@
$L_{n}\mathrel{\mathop{:}}=\lim_{\tau\to\infty}L_{n}(\tau),\quad\text{where}\quad L_{n}(\tau)\mathrel{\mathop{:}}=\int_{0}^{\tau}Y_{n}(r)\mathrm{d}{r}=\int_{0}^{\tau}\lvert\Lambda_{[n]}(r)\rvert\mathrm{d}r.$ (10)
@@
Indeed, the limit exists and is finite for almost all realisations of the ALG, since $(Y_{n}(\tau))_{\tau\geq 0}$ absorbs in 0 almost surely in finite time for any given $n$. In epidemiology, the quantity $L_{n}$, sometimes up to a constant, is referred to as the “cost of an epidemic”, because it measures the total time that individuals spend in the infected state in one episode of the epidemic started with $n$ individuals (Jerwood, 1970; McNeil, 1970; Downton, 1972; Gani and Jerwood, 1972; Ball, 1986; Hernandez-Suarez and Castillo-Chavez, 1999; Andersson and Britton, 2000, Secs. 2.2 and 2.4; Crawford et al., 2018). Previous work has concentrated on its mean $\mathbb{E}[L_{n}]$ or its Laplace-transform $\mathbb{E}[e^{-\theta L_{n}}]$ (or its moment-generating function $\mathbb{E}[e^{\theta L_{n}}]$) but, to the best of our knowledge, explicit expressions for the variance (or higher-order moments) are unavailable so far (but see Stefanov and Wang (2000) for a numerical evaluation of the variance).
2.5 The connection with population genetics
Let us point out the close connections with population genetics theory. The untyped model is closely related with an untyped Moran model with both neutral and selective reproduction events; for review, see Baake and Wakolbinger (2018). There are two main differences: first, cultural traits are not genetically inherited; we therefore do not keep track of the biological parent of a new individual, so the neutral reproduction events of the Moran model turn into death-birth events that only involve a single line. Second, in a learning event, both the incoming and the continuing branches are true cultural parents (recall that learners are counted as their own cultural parents), whereas, in population genetics, they are potential parents, only one of which is the true (genetic) parent.
@@ 2.5 The connection with population genetics
In the same vein, the untyped ALG is closely related to the ASG of Neuhauser and Krone (1997), which contains all potential ancestors of a sample of individuals. More precisely, in our finite-$N$ setting, the untyped ALG is a special case of the killed ASG of Baake et al. (2023), but here without multiple branching, without coalescence events, and without killing. In the typed model, the innovation mechanism is similar to the infinite-sites model of population genetics (see, for example, Wakeley (2009, Ch. 1.2)), where sequences of infinite length are considered and every mutation hits a site that has never mutated before.
@@
If we reduce the number of possible traits to one and do not allow for innovations (that is, consider the limiting case $\mu=0$), we only have two types of individuals: those that have and those that do not have the trait. Let $X(t)\in[N]_{0}$ be the number of individuals with the trait (so $N-X(t)$ do not have it). If $X(t)=x$, we have a transition to $x-1$ at rate $ux$ (death) and to $x+1$ at rate $sx(N-x)/N$ (learning). This is equivalent to a two-type Moran model with selective reproduction at rate $s$ and deleterious mutation at rate $u$, but without neutral and without frequency-dependent reproduction events; that is, a special case of the model tackled by Baake et al. (2023). Furthermore, note that $(X(t))_{t\geq 0}$ has the same birth and death rates as $(Y(\tau))_{\tau\geq 0}$ and hence the same law.
@@
It has been shown by Baake et al. (2023, Theorem 2.3) that the type-frequency process of the Moran model with mutation and frequency-dependent selection on the one hand and the line-counting process of the killed ASG on the other hand are in factorial moment duality with each other. This translates into the factorial moment duality between the type-frequency process of the above single-trait learning model and the line-counting process of the corresponding untyped ALG as follows. For $z,m\in\mathbb{N}_{0}$, let
$z^{\underline{m}}\mathrel{\mathop{:}}=\begin{cases}1,&m=0,\\ z(z-1)\cdots(z-m+1),&m\geq 1\end{cases}$ (11)
@@
Note that $H(x,y)$ is the probability to obtain only individuals without the trait when sampling $y$ individuals without replacement from a population that contains $x$ individuals with the trait (and $N-x$ individuals without it); and the duality means that the sampling probability at time $t$ can be obtained either via the forward or via the backward process. As a consequence, one may obtain, and gain insight into, properties of the learning model forward in time by studying its dual process. See Möhle (1999) for general background on duality in models of population genetics. Since $(X(t))_{t\geq 0}$ and $(Y(\tau))_{\tau\geq 0}$ have the same law, they are actually self dual with respect to $H$.
@@ 2.6 Law of large numbers
In the deterministic limit of the single-trait model, the sequence of processes $(X^{(N)}(t)/N)_{t\geq 0}$ (where the upper index indicates the population size) converges, as $N\to\infty$, to the solution of the initial value problem
@@
provided $X^{(N)}(0)/N\to\xi_{0}$. The differential equation has two equilibria: one at $1-u/s$, the other at 00. They perform a transcritical (or exchange of stability) bifurcation at $s=u$: for $s<u$, the equilibrium at 00 is attracting, while the one at $1-u/s$ (which is then $<0$) is repelling; and vice versa for $s>u$. For $s=u$, the two equilibria coincide at 0, which is then attracting.
The duality (13) carries over to the deterministic limit; specifically, evaluating (13) for $y=1$, $N\to\infty$ such that $x/N\to\xi_{0}$, and $t\to\infty$ gives for the unique stable equilibrium $\bar{\xi}$ of the differential equation that $1-\bar{\xi}=\mathbb{E}[(1-\xi_{0})^{\widetilde{Y}(\infty)}\mid\widetilde{Y}(0)=1]$ for $\xi_{0}\in(0,1]$ and so $\bar{\xi}=1-\mathbb{P}(\lim_{\tau\to\infty}\widetilde{Y}(\tau)=0\mid\widetilde{Y}(0)=1)$. For $s\leq u$, therefore, the trait is always absent at equilibrium, while, for $s>u$, the trait is present in a positive proportion of individuals at the stable equilibrium. See Baake et al. (2018) or Baake and Wakolbinger (2018) for the details about the deterministic limit.
Let us note for later use that the differential equation (14) is not only the deterministic limit of our single-trait model, but also the deterministic limit of $(Y^{(N)}(\tau)/N)_{\tau\geq 0}$ (that is, for $N\to\infty$ in the sequence of processes $(Y^{(N)}(\tau)/N)_{\tau\geq 0}$ with population size $N$, without rescaling of parameters or time), provided that $Y^{(N)}(0)/N$ converges; this is clear because $(X(t))_{t\geq 0}$ and $(Y(t))_{t\geq 0}$ have the same law.
In the following, we continue to adhere to finite $N$, but the two regimes $s<u$ and $s>u$ still behave in qualitatively different ways (as long as $s$ is not too close to $u$ and with a smooth transition between the two regimes), see Nåsell (2011, Chap. 2.5) or Foxall (2021). We will therefore continue to use the notions subcritical and supercritical, in line with the literature.

### v2

@@ 2.4 The ALG as a stochastic process
So far, we have described the ALG as a function of a given realisation of the graphical representation. We now define it as a stochastic process independent of such a realisation, again first in the untyped and then in the typed version.
@@
    For the untyped version, let $\Lambda^{t}_{\alpha}(\tau)\subseteq[N]$ denote the set of ancestors at backward time $\tau\geq 0$ of individual $\alpha\in[N]$ at forward time $t\geq 0$; note that we allow $\tau>t$ as before. For $G\subseteq[N]$, set $\Lambda^{t}_{G}(\tau)\mathrel{\mathop{:}}=\cup_{\alpha\in G}\Lambda^{t}_{\alpha}(\tau)$ by slight abuse of notation (so $\Lambda^{t}_{\{\alpha\}}(\tau)=\Lambda^{t}_{\alpha}(\tau)$). Then the process $(\Lambda(\tau))^{t}_{\tau\geq 0}\mathrel{\mathop{:}}=\big(\Lambda^{t}_{1}(\tau),\ldots,\Lambda^{t}_{N}(\tau)\big)_{\tau\geq 0}$ contains the complete genealogical information of all individuals in the present population. It evolves as follows. Clearly $(\Lambda^{t}_{1}(0),\ldots,\Lambda^{t}_{N}(0))=(\{1\},\{2\},\ldots,\{N\})$. If, for any time $\tau\geq 0$, the current state is $\Lambda^{t}(\tau)=(\lambda_{1},\ldots,\lambda_{N})$, the following events may happen: at rate $Nu$, a death-birth event hits a randomly-chosen individual $\beta$. So $\beta$ is removed from the set of ancestors of anybody, and
@@
    At rate $Ns$, a learning event occurs and a randomly-chosen individual $\beta_{1}$ learns from another randomly-chosen individual $\beta_{2}$ (where $\beta_{2}=\beta_{1}$ is allowed, in which case the event is silent). So if, for $\alpha\in[N]$, $\beta_{1}$ is contained in $\lambda_{\alpha}$, then $\beta_{2}$ joins this ancestral set, and otherwise nothing happens:
    $\lambda_{\alpha}\longrightarrow\begin{cases}\lambda_{\alpha}\cup\{\beta_{2}\},&\text{if }\beta_{1}\in\lambda_{\alpha},\\ \lambda_{\alpha},&\text{otherwise},\end{cases}\qquad\alpha\in[N].$ (6b)
@@
    The untyped ALG may be turned into a typed one by superposing the untyped graph with a Poisson point process that lays down innovations at rate $\mu$ on every line segment. Together with an initial assignment of types to the lines alive at backward time $t$ (recall this step is void if we run the graph until there are no lines left) and propagation along the graph as in the forward model, this determines the types along all lines in the graph.
Returning to the untyped version, let us note for later use that, for any $n\in[N]$ and $\alpha\in[n]$, we have $\Lambda^{t}_{\alpha}(\tau)\subseteq\Lambda^{t}_{[n]}(\tau)$ for all $t,\tau\geq 0$ and so
$|\Lambda^{t}_{\alpha}(\tau_{1})|=|\Lambda^{t}_{[n]}(\tau_{1})|\quad\text{for some $\tau_{1}\geq 0$ implies}\quad\Lambda^{t}_{\alpha}(\tau)=\Lambda^{t}_{[n]}(\tau)\hskip 10.0pt\text{for all }\,\tau\geq\tau_{1}.$ (7)
In what follows, an important role will be played by the line-counting process $(Y^{t}_{n}(\tau))_{\tau\geq 0}$ of the ALG. That is,
$Y^{t}_{n}(\tau)\mathrel{\mathop{:}}=\lvert\Lambda^{t}_{[n]}(\tau)\rvert$ (8)
is the number of lines at backward time $\tau$ of the sample $[n]$ of individuals taken at forward time $t$; due to exchangeability, we allow ourselves to simply speak of a sample of size $n$. Also, we will sometimes omit the dependence on $n$. For any given $t$, $(Y^{t}(\tau))_{\tau\geq 0}$ is a birth-death process [^1] This birth-death process is not to be confused with the death-birth events in the model, as indicated by crosses in the graphical representation. on $N]_{0}:=\{0,1,\ldots,N\}$ with initial value $Y^{t}_{n}(0)=n$ and birth and death rates222not to be confused with the realisations $\lambda_{\alpha}$ of $\Lambda_{\alpha}$ and the innovation rate $\mu$.
@@
respectively, when $Y^{t}(\tau)=y$ (note that the factor (Ny)/N(N-y)/N is the probability that a learning arrow comes from outside the current graph). Clearly, $(Y^{t}(\tau))_{\tau\geq 0}$ is an absorbing Markov chain with 0 as the only absorbing state. The behaviour of the process is very well studied, since it is, at the same time, the stochastic logistic process of ecology and epidemiology; in particular, $(Y^{t}_{n}(\tau))_{\tau\geq 0}$ is the number of infected individuals in a stochastic SIS model with infection rate $s$ and recovery rate $u$ per individual when started with $n$ infected individuals at time $t$ (Andersson and Britton, 2000, Ch. 8.2).
@@
As alluded to already, another crucial quantity will be $L^{t}_{n}$, the (total) length of the genealogy, that is, the total length of all branches, as $\tau\to\infty$, in the ALG of a sample of size $n$ taken at time $t$, that is,
@@
$L^{t}_{n}\mathrel{\mathop{:}}=\lim_{\tau\to\infty}L^{t}_{n}(\tau),\quad\text{where}\quad L^{t}_{n}(\tau)\mathrel{\mathop{:}}=\int_{0}^{\tau}Y^{t}_{n}(r)\mathrm{d}{r}=\int_{0}^{\tau}\lvert\Lambda^{t}_{[n]}(r)\rvert\mathrm{d}r=\int_{t-\tau}^{t}\lvert\Lambda^{t}_{[n]}(t-s)\rvert\mathrm{d}s.$ (10)
@@
Indeed, the limit exists and is finite for almost all realisations of the ALG, since $(Y^{t}_{n}(\tau))_{\tau\geq 0}$ absorbs in 0 almost surely in finite time for any given $n$. In epidemiology, the quantity $L^{t}_{n}$, sometimes up to a constant, is referred to as the cost of an epidemic, because it measures the total time that individuals spend in the infected state in one episode of the epidemic started with $n$ individuals at time $t$ (Jerwood, 1970; McNeil, 1970; Downton, 1972; Gani and Jerwood, 1972; Ball, 1986; Hernandez-Suarez and Castillo-Chavez, 1999; Andersson and Britton, 2000, Secs. 2.2 and 2.4; Crawford et al., 2018). Previous work has concentrated on its mean $\mathbb{E}[L^{t}_{n}]$ or its Laplace transform $\mathbb{E}[e^{-\theta L^{t}_{n}}]$ (or its moment-generating function $\mathbb{E}[e^{\theta L^{t}_{n}}]$) but, to the best of our knowledge, explicit expressions for the variance (or higher-order moments) are unavailable so far (but see Stefanov and Wang (2000) for a numerical evaluation of the variance).
Let us emphasise that we have defined the ancestral processes as functions of the forward time $t$ where we start looking back; this will become important in Section 3.4.2. For the moment, we fix $t=t_{\text{max}}$ and, when we do so, omit the superscript, so $\Lambda_{\alpha}(\tau)\mathrel{\mathop{:}}=\Lambda^{t_{\text{max}}}_{\alpha}(\tau)$ and likewise for the other quantities.
2.5 The connection with models of population genetics
We have already mentioned the connection of the ALG with the SIS model and, in particular, the stochastic logistic process; this will be used extensively later. The genealogical point of view, however, is less important in epidemiology; notably, in epidemiology, the stochastic logistic process usually appears in forward-time models, whereas our $(Y(\tau))_{\tau\geq 0}$ is an ancestral process. It is therefore worthwhile to investigate the somewhat subtle connection with population genetics theory more closely, where the backward perspective is crucial and highly developed.
We claim that our learning model is closely related to, but not identical with, a Moran model with selective reproduction and deleterious mutation, but without resampling (see Baake and Wakolbinger (2018) for a review of the Moran model and its ancestral structures under the various evolutionary forces mentioned). To see this, let us first consider the Moran model with selective reproduction, deleterious mutation, and with resampling, and then compare it to the learning model. In this model, selective arrows appear between every ordered pair of lines at rate $s/N$; both the incoming and the continuing branches are then potential genetic parents of the affected individual, with only one of them being the true parent. These arrows correspond to the learning arrows of our model. The difference is that both the incoming and the continuing lines are (true) cultural parents (recall that learners are counted as their own cultural parents); there are no potential parents. In line with this, a learner’s knowledge set is not replaced by, but merges with that of the cultural parent. Furthermore, it is important to note that, while the selective arrows in the Moran model represent (genetic) selection, the corresponding learning arrows in the ALG are (culturally) neutral since individuals are chosen as cultural parents independently of the traits they carry. Next, resampling is a central feature of the Moran model and means that selectively neutral reproduction arrows appear at rate $1/N$ between every ordered pair of lines. They indicate that the individual at the tail is the genetic parent of the newborn individual at the tip of the arrow. Since resampling is absent in the learning model, we speak of the latter as a Moran-type model. Finally, deleterious mutation events in the Moran model happen at rate $u$ per line and mean that the type of that individual is ‘reset’ to the (selectively) weakest type. The corresponding events in the learning model are the death-birth events, where an individual is replaced by an ignorant newborn.
The connection is mirrored in the corresponding ancestral structures. Indeed, the untyped ALG is identical with a finite-population variant of the untyped ASG that lacks coalescence events (because there is no resampling), and where a line is pruned if it is hit by a deleterious mutation. The difference lies in the interpretation: the lines of the ALG correspond to the cultural parents of the sample, whereas the lines of the pruned ASG represent the potential genetic parents.
As a consequence, the ALG belongs into the context of the rich recent literature on variants of the ASG and their applications. For example, Pfaffelhuber and Pokalyuk (2013) and Greven et al. (2016) used the ASG to study fixation of a beneficial mutant; Cordero (2017) studied the ASG in a law-of-large-numbers regime (note that here, as well, the coalescence events vanish); Cordero and Möhle (2019) studied the stationary number of lines under generalised coalescence mechanisms; Boenkost et al. (2021) extended the ASG to the Cannings model of reproduction; Cordero and Véchambre (2023) investigated it in a random environment; and Baake et al. (2023) included multiple branching, pruning individual lines, and killing the entire process in order to account for a certain kind of frequency-dependent selection, deleterious, as well as beneficial mutations.
@@ 2.5 The connection with models of population genetics
In the typed learning model, the innovation mechanism is similar to the mutation process in the infinite-sites model of population genetics (see, for example, Wakeley (2009, Ch. 1.2)), where sequences of infinite length are considered and every mutation hits a site that has never mutated before.
@@
If we reduce the number of possible traits to one and do not allow for innovations (that is, consider the limiting case $\mu=0$), we only have two types of individuals: those that have and those that do not have the trait. Let $X(t)\in[N]_{0}$ be the number of individuals with the trait (so NX(t)N-X(t) do not have it). If $X(t)=x$, we have a transition to x1x-1 at rate $ux$ (death) and to $x+1$ at rate sx(Nx)/Nsx(N-x)/N (learning). This is equivalent to a two-type Moran model with selective reproduction at rate $s$ and deleterious mutation at rate $u$, but without resampling, without frequency-dependent selection, and without beneficial mutations; it is a special case of the model tackled by Baake et al. (2023). Furthermore, note that $(X(t))_{t\geq 0}$ has the same birth and death rates as $(Y(\tau))_{\tau\geq 0}$ and hence the same law.
@@
The type-frequency process of the Moran model with resampling, frequency-dependent selection, as well as deleterious and beneficial mutation on the one hand and the line-counting process of the corresponding ASG variant on the other hand are in factorial moment duality with each other (Baake et al., 2023, Theorem 2.3). This translates into the factorial moment duality between the type-frequency process of the above single-trait learning model and the line-counting process of the corresponding untyped ALG as follows. For $z,m\in\mathbb{N}_{0}$, let
$z^{\underline{m}}\mathrel{\mathop{:}}=\begin{cases}1,&m=0,\\ z(z-1)\cdots(z-m+1),&m\geq 1,\end{cases}$ (11)
@@
Note that $H(x,y)$ is the probability to obtain only individuals without the trait when sampling $y$ individuals without replacement from a population that contains $x$ individuals with the trait (and NxN-x individuals without it); and the duality means that the sampling probability at time $t$ can be obtained either via the forward or via the backward process. As a consequence, one may obtain, and gain insight into, properties of the learning model forward in time by studying its dual process. Since $(X(t))_{t\geq 0}$ and $(Y(\tau))_{\tau\geq 0}$ have the same law, they are actually self dual with respect to $H$.
Let us mention that factorial moment dualities have a long history in the context of fixed-size population genetic models. In the case without selection, they already appear in papers by Cannings (1974) and Gladstien (1977a, b, 1978), at a time where neither genealogical processes nor the concept of dualities for Markov chains had been formulated yet. Rather, the dualities appear in terms of algebraic identities between matrices, and are used to calculate the eigenvalues of the Markov transition matrices via a similarity transform; the connection with the backward point of view is at most implicit. In 1999, Möhle established factorial moment dualities in neutral genetic models, explicitly and in terms of backward processes; this paper also provides general background on duality in models of population genetics.
@@ 2.6 Law of large numbers
In the deterministic limit of the single-trait model $(X(t))_{t\geq 0}$ as described in the previous section, the sequence of processes $(X^{(N)}(t)/N)_{t\geq 0}$ (where the upper index indicates the population size) converges, as $N\to\infty$, to the solution of the initial value problem
@@
provided $X^{(N)}(0)/N\to\xi_{0}$. The differential equation has two equilibria: one at 1u/s1-u/s, the other at 0. They perform a transcritical (or exchange of stability) bifurcation at $s=u$: for $s<u$, the equilibrium at 0 is attracting, while the one at 1u/s1-u/s (which is then <0<0) is repelling; and vice versa for $s>u$, where 1u/s>01-u/s>0. For $s=u$, the two equilibria coincide at 0, which is then attracting.
The duality (13) carries over to the deterministic limit; specifically, evaluating (13) for $y=1$, $N\to\infty$ such that $x/N\to\xi_{0}$, and $t\to\infty$ gives for the unique stable equilibrium $\bar{\xi}$ of the differential equation that $1-\bar{\xi}=\mathbb{E}[(1-\xi_{0})^{\widetilde{Y}(\infty)}\mid\widetilde{Y}(0)=1]$ for $\xi_{0}\in(0,1]$ and so $\bar{\xi}=1-\mathbb{P}(\lim_{\tau\to\infty}\widetilde{Y}(\tau)=0\mid\widetilde{Y}(0)=1)$. For $s\leq u$, therefore, the trait is always absent at equilibrium, while, for $s>u$, the trait is present in a positive proportion of individuals at the stable equilibrium. See Cordero (2017), Baake et al. (2018), or Baake and Wakolbinger (2018) for the details about the deterministic limit.
Let us note for later use that the differential equation (14) is not only the deterministic limit of our single-trait model, but also the deterministic limit of $(Y^{(N)}(\tau)/N)_{\tau\geq 0}$ (that is, for $N\to\infty$ in the sequence of processes $(Y^{(N)}(\tau)/N)_{\tau\geq 0}$ with population size $N$, without rescaling of parameters or time), provided that $Y^{(N)}(0)/N$ converges; this is clear because $(X^{(N)}(t))_{t\geq 0}$ and $(Y^{(N)}(\tau))_{\tau\geq 0}$ have the same law.
In what follows, we continue to adhere to finite $N$, but the two regimes $s<u$ and $s>u$ still behave in qualitatively different ways (as long as $s$ is not too close to $u$ and with a smooth transition between the two regimes), see Nåsell (2011, Chap. 2.5) or Foxall (2021). We will therefore continue to use the notions subcritical and supercritical, in line with the literature.